[Hallc_running] [Revised Logentry] RC Update

smithg at jlab.org smithg at jlab.org
Fri Nov 22 18:15:02 EST 2019

Logentry Text:
I thought it would be useful to add an update describing some of what happened on a very busy day:

This morning we gathered a small army to do the fiber pull. Joe had set up this flexible orange conduit from the Hall A counting room, under the floor to the Hall C counting room, then to the cable shaft in the Hall C counting room, along the hall from the key room to the door entering Hall C, up 3 stories to the cryoline tray, then over to the cryotarget platform, then to the laser box. Heroic efforts from Arun, Murchana, Junhao, Mingyu, Melanie, Larry, Rob, Chuck, and of course JOE. Hope I didn't forget anyone. It was really difficult. The bundle kept getting stuck. We had to pull with all our might and  sometimes go back and forth to get it unstuck, which required a lot of running back and forth. These fibers are 110 m long! Needed a JLG in the end to violently shake the conduit up near the cryo cable tray to get a few inches at a time pulled through. One time it got stuck it was due to a single fiber in a small diameter knot (kiss that fiber goodbye). But it's pulled and!
  there are a couple meters in the laser box... Don't forget to put the cover back on the laser box this weekend.

During this time the laser hut inspection got punted to Monday at 1:30 PM, for a number of reasons.

The good news is that the US beamline spool piece with the Be window was installed, aligned, and the US beamline pumps are on, so the US girder can be aligned early next week. In fact, let me just cut & paste from  Chris Gould's email summary of the survey group's work:

We completed the ladder survey with Ethan's support.  We decided to leave it in the carbon hole position thinking that this would help Arun with his work this weekend.  The laser fiducials are also installed and set coincident with beam.

The Be spool piece is aligned.  Our plan is to survey the diagnostic girder on Monday.  Please let me know if this conflicts with other activities.

I believe the only surveys remaining are the diagnostic girder and pointing surveys, correct me if I am wrong.  JP, let me know if you decided whether or not you need support with the optics table.

OK, now back to me (RC Greg). I think that the plan this weekend is for the target people to start working with the laser alignment, now that it's all in place, using just a simple pocket laser pointer though for now. We're making a lot of progress. Lots left to do though! The next potential problem is the downstream beamline- we still don't have the Be window spool piece for that. A priority for next week.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3743342
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20191122/ecfbf8a5/attachment.html>

More information about the Hallc_running mailing list